# Decoding Open-Ended Information Seeking Goals from Eye Movements in Reading

### v1

@@ 7 Related Work
Most notably, two studies have addressed decoding tasks for reading goals. Hollenstein et al. (2021b) classified whether participants were engaged in ordinary reading or annotating semantic relations in single sentences using the ZuCo corpus Hollenstein et al. (2020, 2023). Closest to our study is Shubi et al. (2024a), who classified ordinary reading versus information seeking using the OneStop dataset. In both cases, the decoding tasks are procedural – classifying between two types of reading. In contrast, our study addresses a related but conceptually and practically different task which is not procedural, but rather semantic and text specific – decoding reading tasks specific to the text, within the information seeking regime. In our setup, instead of a small number of categories that can apply to any text, we have hundreds of text-specific tasks that can be arbitrary in nature.
@@ 8 Summary and Discussion
In this work, we introduced a new task: decoding of arbitrary information seeking goals over specific texts from eye movements in reading. This task has both scientific value and practical potential, as decoding of information seeking goals in real time can open the door for real-world applications in various domains, including education, e-learning and content accessibility. We tackle this task in two formulations using a number of modeling approaches both existing and new. Overall, we find that despite the inherent noise in eye movement data, highly valuable information can be extracted from such data regarding specific information seeking goals. While the presented results are highly encouraging, ample room remains for further improvement of prediction accuracy via modeling innovations.
9 Limitations
Our study relies on the OneStop dataset. While OneStop is uniquely suited for addressing the proposed task, our work has a number of limitations stemming from the experimental setup and the eye movement recordings in this dataset. The information-seeking tasks in OneStop are at the granularity level of single paragraph, leaving out information seeking in both shorter and longer texts. The questions further adhere to the STARC annotation framework Berzak et al. (2020) which does not include tasks that require making inferences over the entire text, nor verbatim string matching tasks, which are both common in daily life. The texts themselves cover a wide range of topics, but are restricted to the newswire domain. New experimental designs and additional data collection work are needed to address these limitations.
An additional limitation of the current study is the restriction to native English speakers. Data from the most proficient group of readers for a given language may lead to overestimates of performance in the general population, which includes lower skill readers such as L2 speakers, children, participants with cognitive impairments and others. Furthermore, the study is restricted to English, and additional data collection and modeling work is needed for other languages. We plan to collect such data in future work.
A further general drawback of in-lab eye tracking studies is that the lab setting may lead to reading patterns that deviate from everyday life reading (Huettig and Ferreira, 2022). Moreover, the eye tracking experiment is conducted using a high-end Eyelink 1000 Plus eye tracker at a sampling rate of 1000Hz. To increase the relevance of our work for deployment in real-world scenarios, such as web browsing, additional studies are required to test the feasibility of the proposed tasks and methods when the eye tracking data is collected using equipment with a lower sampling rate, lower tracking precision or both.
@@ 10 Ethical Considerations
The eye movements dataset used in this work was collected by Berzak et al. (2025). The study was conducted under an institutional IRB protocol, and all the participants provided written consent before participation. The data is anonymized. Analyses and predictive modeling of task-based reading are among the primary use cases for which the data was collected.
While the current work is a scientific proof of concept and is not aimed at a particular user-facing application, one has to consider potential use cases and risks for the presented task. Automatic decoding of reading goals can have multiple potential applications that offer societal benefits. It can be used as part of e-tutoring systems that assess students’ progress on various types of information-seeking tasks and monitor their information-seeking skills over time. An ability to automatically determine specific information needs in real-time can also be used to assist special populations like elderly people when accessing information on the web. However, it is important to note that such technologies, especially in their current level of accuracy, may be unreliable and introduce biases for various individuals and groups, such as L2 readers and groups with reading and cognitive impairments. Additional data collection and analyses are needed to assess such biases.
@@
Appendix A Model Training and Hyperparameters
@@ A.1 Discriminative Models
For all neural models, we train with learning rates of $\{0.00001,0.00003,0.0001,0.0002\}$, following Shubi et al. (2024b). Additionally, for all models that make use of word embeddings, we include both frozen and unfrozen language model variants in the search space.
  • For Haller RNN, we search over LSTM layer counts of $\{1\}$, hidden sizes of $\{10,40,70,140\}$, and a dropout rate of $0.1$. The model is trained with and without freezing language model parameters.
@@
We train the deep-learning-based models for a maximum of 40 epochs, with early stopping after 8 epochs if no improvement in the validation error is observed. A single training epoch took roughly 5 minutes for RoBERTa-F, 10 minutes for Haller RNN, and 30 minutes for Llava and Llama. Each individual run was capped at 24 hours. Following Liu et al. (2019); Mosbach et al. (2021); Shubi et al. (2024b), we use the AdamW optimizer (Loshchilov and Hutter, 2018) with a batch size of $16$. RoBERTa-F uses a linear warm-up ratio of $0.06$ and a weight decay of $0.1$. We standardize each eye movement feature using statistics computed on the training set, to zero mean unit variance.
Both classification baseline models use word embeddings from the RoBERTa-Large (Liu et al., 2019) language model.
A.2 Generative Models
@@
This model employs a teacher-forcing strategy. Specifically, loss computation is restricted exclusively to tokens corresponding to possible questions, contrasting the likelihood of generating the correct question against two distractors. The computed losses for each question candidate are inverted, serving as logits for softmax normalization and subsequently optimized via cross-entropy against the true label. The LLaVA backbone remains frozen, and training employs LoRA with dropout rate of 0.1, RS-LoRA, and rank $r=8$. The hyperparameter search includes learning rates of ${1\times 10^{-5},3\times 10^{-5},1\times 10^{-4}}$, fixation embedding hidden sizes of ${512,1024}$, and early stopping after 8 epochs without validation improvement.
All neural network-based models were trained using the PyTorch Lightning (Falcon and team, 2024) library on NVIDIA A100-40GB and L40S-48GB GPUs.
We utilize the Hugging Face implementation of LLaVA-OneVision, specifically LLaVA-hf/LLaVA-onevision-qwen2-0.5b-si-hf. Additionally, we employ the gpt-4o-2024-08-06 version of GPT-4o. We use Unsloth (Daniel Han and team, 2023) to fine-tune the Llama model. For the QA model we use LIAMF-USP/roberta-large-finetuned-race.

### v2

@@ 7.2 Decoding Cognitive State from Eye Movements
Most notably, two studies addressed decoding of the reader’s goals. Hollenstein et al. (2021b) classified ordinary reading for comprehension versus annotation of semantic relations in single sentences using the ZuCo corpus (Hollenstein et al., 2020; 2023). Shubi et al. (2025) classified ordinary reading versus information seeking using the OneStop dataset. Importantly, in both studies, the decoding tasks are procedural, where the goal is to distinguish between a small set of pre-defined manners of reading that are not specific to a particular text. In contrast, our study addresses a related but conceptually and practically different task, which is not procedural, but rather semantic, decoding reading tasks specific to the text, within the information seeking regime. Instead of a small number of categories that can apply to any text, we have hundreds of text-specific tasks that can be arbitrary in nature. See further discussion on the procedural versus semantic tasks distinction in Shubi & Berzak (2023). Furthermore, these and other studies on cognitive state decoding typically use discriminative, encoder-based models that are limited to classification tasks. Here, we also integrate eye movements into generative, decoder-based language models, and investigate both textual- and embedding-based methods for conditioning text generation on eye movements in reading.
@@ 8 Summary and Discussion
We introduce a new challenge of both scientific and practical value: decoding of arbitrary information seeking goals over specific texts from eye movements in reading. We tackle this challenge in two formulations, goal selection and goal reconstruction, using a number of evaluation frameworks and modeling approaches. The best performing model on the selection task is able to predict the correct question with a considerable degree of success, even when the candidate reading goals pertain to the same information in the text. Our results further suggest that although the reconstruction task is extremely challenging, meaningful progress can also be made in this open-ended regime. Overall, we find that despite the inherent noise in eye movement data, effective modeling can extract highly valuable information regarding specific information seeking goals at the challenging granularity level of eye movements of a single reader over a single paragraph.
Automatic decoding of information seeking goals in real time can pave the way for new applications with positive societal value. In the future, media outlets and providers of municipal and governmental services could better understand the information needs of users who access their websites and render information in a manner that better meets these needs. Automatic identification of information seeking goals can also be used for real-time assistance to special populations, like elderly people, when accessing critical information on the web. Future e-learning systems could assess students’ progress on various types of information-seeking tasks and monitor their information-seeking skills over time. Finally, one can imagine interactive reading interfaces that include real-time text personalization, such as simplification and suggestion of additional relevant information, according to the reader’s information seeking goals. The present study is a first step in enabling such applications. Ample room remains for further improvement of prediction accuracy via modeling innovations, as well as the collection and analysis of data for additional variants of information seeking regimes, corpora, reader populations, and languages.
@@ 9 Ethical Considerations
The OneStop Eye Movements dataset used in this work was collected by Berzak et al. (2025). The study was conducted under an institutional IRB protocol, and all the participants provided written consent before participation. The data is anonymized. Analyses and predictive modeling of task-based reading are among the primary use cases for which the data was collected.
While the current work is a scientific proof of concept and is not aimed at a particular user-facing application, one has to consider potential use cases and risks for the presented task. In particular, it is important to note that given the current level of accuracy of goal decoding, educational and assistive technologies that rely on this task may be unreliable and introduce biases for various individuals and groups, such as L2 readers and groups with reading and cognitive impairments. Additional data collection and analyses are needed to assess such biases.
@@
10 Reproducibility Statement
We describe the model training and selection procedure, the evaluation protocol, the hyperparameter search space for each model, and the hardware and software specifications in Sections˜5 and C. The OneStop dataset (Berzak et al., 2025) used in the experiments and described in Section˜2 is publicly available at https://osf.io/2prdq/. The code for the paper, which implements all the models, experimental procedures, and analyses, is available at: https://anonymous.4open.science/r/open-question-prediction/.
@@ 10 Reproducibility Statement
Appendix A Discriminative Models
The figures and descriptions below, depicting model architectures, show the processing of a single candidate question. As mentioned in Section˜4, each candidate question is processed independently.
A.1 Haller RNN
Given a text (a paragraph in our case) $T$ and a candidate question $q\in Q_{T}$, the model suggested by Haller et al. (2022) and adapted by us performs the following steps:
  1. 1.
    Word Embeddings: LLM-based contextualized word embeddings, for $T$ and $q$, resulting in three subsequences: (1) $Z_{[CLS]}$, (2) $Z_{q_{t_{1}}},\dots,Z_{q_{t_{l}}}]$, and (3) $Z_{T_{t_{1}}},\dots,Z_{T_{t_{x}}}]$.
  2. 2.
    Fixation Sequence: Paragraph-only token-level embeddings are pooled to word-level embeddings, and ordered by the fixation sequence. Then, each embedding in the sequence is concatenated with the respective fixation features. The fixation-level features are:
    • Horizontal position of the fixation
    • Total gaze duration (sum of all fixations on the word)
    • Duration of first fixation
    • Duration of outgoing saccade
    • Horizontal distance of outgoing saccade
    • Vertical distance of outgoing saccade
    • Total distance of outgoing saccade
    • Duration of incoming saccade
    • Horizontal distance of incoming saccade
    • Vertical distance of incoming saccade
  3. 3.
    Projection: $Z_{[CLS]}$ and $Z_{q_{t_{1}}},\dots,Z_{q_{t_{l}}}]$ are projected to a higher dimension to fit the dimension of the concatenated embeddings from the previous step.
  4. 4.
  5. 5.
Figure˜5(a) depicts the architecture adapted to our task.
A.2 RoBERTEye-Fixations
Formally, given a paragraph $T$ and a candidate question $q\in Q_{T}$, the model constructs two parallel representations:
  1. 1.
    Textual Representation ($Z_{W}$): The input follows the format $Z_{W}=[\texttt{CLS};T;\texttt{SEP};q;\texttt{SEP}]$.
  2. 2.
    Fixation-Level Eye Movement Representation ($Z_{E_{T}}$): The eye movement sequence is defined as $Z_{E_{T}}=[Z_{E_{f_{1}}},...,Z_{E_{f_{m}}}]$. The fixation representation is computed as:
    Here, $E_{f_{i}}$ captures fixation properties (e.g., duration, position) for fixation $f_{i}$ on word $w_{i}$. The fully connected layer FC projects this feature vector into the word embedding space, $\text{Emb}_{\text{pos}}(w_{i})$ is the positional embedding of $w_{i}$ used to map each fixation to the corresponding word, and $\text{Emb}_{\text{eye}}$ is a learnable embedding marking the presence of eye movement information. The fixation-level features are described in Table˜3.
  3. 3.
    Fusion and Prediction: The combined sequence $Z_{E_{T}};\texttt{SEP}_{E};Z_{W}]$ is processed by the transformer encoder, and the final CLS token representation is passed through two fully connected layers to predict the question: $\hat{q}=\operatorname{argmax}\;\text{MLP}(\texttt{CLS})$.
Figure˜5(b) depicts the adapted architecture.
Table 3: Eye movement and word property features used by RoBERTEye-Fixations and DalEye-Llama. See Berzak et al. (2025) for further details.
Feature Name Description
Word-Level Eye Movement Features
IA_DWELL_TIME The sum of the duration across all fixations that fell in the current interest area
IA_DWELL_TIME_% Percentage of trial time spent on the current interest area.
IA_FIXATION_% Percentage of all fixations in a trial falling in the current interest area.
IA_FIXATION_COUNT Total number of fixations falling in the interest area.
IA_REGRESSION_IN_COUNT Number of times interest area was entered from a higher IA_ID.
IA_REGRESSION_OUT_FULL_COUNT Number of times interest area was exited to a lower IA_ID.
IA_RUN_COUNT Number of times the Interest Area was entered and left (runs).
IA_FIRST_FIX_PROGRESSIVE Checks whether the first fixation in the interest area is a first-pass fixation.
IA_FIRST_FIXATION_DURATION Duration of the first fixation event that was within the current interest area
IA_FIRST_FIXATION_VISITED_IA_COUNT Number of distinct interest areas visited before the first fixation on the current area.
IA_FIRST_RUN_DWELL_TIME Dwell time of the first run.
IA_FIRST_RUN_FIXATION_COUNT Number of all fixations in a trial falling in the first run of the current interest area.
IA_SKIP IA_SKIP = 1 if the area received no fixation in first-pass reading.
IA_REGRESSION_PATH_DURATION Total fixation duration on the current interest area until moving to a higher IA_ID.
IA_REGRESSION_OUT_COUNT Exits from current IA to lower IDs before fixating a higher ID.
IA_SELECTIVE_REGRESSION_
PATH_DURATION
Duration of fixations and refixations on the current
interest area until eyes move to a higher ID.
IA_LAST_FIXATION_DURATION Duration of the last fixation event that was within the current interest area.
IA_LAST_RUN_DWELL_TIME Dwell time of the last run.
IA_LAST_RUN_FIXATION_COUNT Number of fixations during the last run (sequence) within the current interest area.
IA_TOP Y coordinate of the top of the interest area.
IA_LEFT X coordinate of the left-most part of the interest area.
IA_FIRST_FIX_PROGRESSIVE Whether the first fixation in the interest area is progressive (left-to-right, first-pass).
normalized_Word_ID Position in the paragraph of the word interest area, normalized from zero to one.
PARAGRAPH_RT Reading time of the entire paragraph.
total_skip Binary indicator whether the word was fixated on.
Fixation-level Eye Movement Features
CURRENT_FIX_INDEX The position of the current fixation in the trial.
CURRENT_FIX_DURATION Duration of the current fixation.
CURRENT_FIX_PUPIL Average pupil size during the current fixation.
CURRENT_FIX_X X coordinate of the current fixation.
CURRENT_FIX_Y Y coordinate of the current fixation.
NEXT_FIX_ANGLE, PREVIOUS_FIX_ANGLE Angle between the horizontal and the line from the current to the next/previous fixation.
NEXT/PREVIOUS_FIX_DISTANCE Distance between current fixation and next/previous fixation.
NEXT_SAC_AMPLITUDE Amplitude of the following saccade in degrees of visual angle.
NEXT_SAC_ANGLE Angle between the horizontal plane and the direction of the next saccade.
NEXT_SAC_AVG_VELOCITY Average velocity of the next saccade.
NEXT_SAC_DURATION Duration of the next saccade in milliseconds.
NEXT_SAC_PEAK_VELOCITY Peak values of gaze velocity (in visual degrees per second) of the next saccade.
NEXT_FIX_INTEREST_AREA_INDEX Index of the interest area where the next fixation lands
Word Properties
gpt2_surprisal Difficulty predicted by GPT-2 model
wordfreq_frequency Frequency of word in language usage
word_length Number of characters in the word
start_of_line Indicates word is at start of line
end_of_line Indicates word is at end of line
is_content_word Is the word a content word (noun, verb, etc.)
left_dependents_count Count of dependents left of word
right_dependents_count Count of dependents right of word
distance_to_head Syntactic distance to head word
(a) Haller RNN
(b) RoBERTEye-Fixations
Figure 5: Visualization of the Haller RNN and RoBERTEye-Fixations architectures. LMLM stands for a language model, RNNRNN for a recurrent neural network. $FF_{f_{i}}$ stands for the fixation features and $w_{f_{i}}$ for the word corresponding to the $i$-th fixation respectively.
Appendix B Generative Models
B.1 DalEye-LLaVA
DalEye-LLaVA extends the LLaVA-OneVision architecture by replacing the vision encoder with a fixation encoder tailored to eye movement features. The model input is structured as an instruction-style prompt comprising three components: (1) a task description, (2) the paragraph text, and (3) the participant’s eye movement features, positionally aligned with the paragraph words. The prompt concludes with the ground-truth question followed by an <eos> token. The entire sequence is tokenized, and training uses teacher forcing, with the cross-entropy loss applied only to the question tokens.
The forward architecture mirrors LLaVA’s design, with three main stages, as depicted in Figure˜6:
  1. 1.
    Text encoding: The task prompt and paragraph tokens are embedded by the backbone language model.
  2. 2.
    Fixation encoding: Eye movement features (see Table˜3) are processed by a dedicated fixation encoder, which applies MLP layers across the feature dimension and 1D convolutions across the temporal dimension, producing a sequence of fixation embeddings.
  3. 3.
    Fusion: Fixation embeddings replace the <image> placeholder token in the input sequence. Word tokens are assigned positional IDs according to their order in the paragraph, while each fixation embedding inherits the positional ID of the corresponding fixated word.
Figure 6: DalEye-LLaVA architecture. Text tokens and eye movement features are fused via a fixation encoder and rotary positional embeddings before being passed to the LLM.
B.2 DalEye-Llama
In Section˜B.2.1 we detail the task prompt and eye-movement input representations used for question generation with DalEye-Llama. Variants of these prompts and inputs are described in Section˜B.2.2.
B.2.1 Input format
You will be given data from an eye-tracking for reading experiment in which participants first read a question about a paragraph, then read the paragraph, and finally answered the question.
Input: You will be provided with a paragraph and eye movements of a single participant over that paragraph.
Output: Your task is to generate the question that was presented to the participant prior to reading the paragraph.
Eye Movements Representation:
You will receive the eye movements data for the paragraph formatted as <FORMAT>
Instructions:
Output only the original question presented to the reader, matching it as best as you can. DO NOT include any additional commentary or explanation.
<PARAGRAPH>
<SCANPATH>
<PARAGRAPH> is the textual paragraph, for example: ‘‘The quick brown fox jumps...’’.
<SCANPATH> is the eye movements data which depends on the <FORMAT>.
B.2.2 Input and Task Prompt Variations
We experiment with three input representations that capture different granularities of eye movement information: (i) Fixation—level, (ii) Word-level, and (iii) a combination of both fixation- and word-level information, as detailed below:
  1. 1.
    Fixations: Fixation Duration + Next saccade direction
    • Format: a list of fixation-level features: [fixated word index, fixated word, fixation duration in ms, direction of next saccade (backward to earlier word / within word / forward to later word), words between current and next fixation]
    • Example: [[4, ‘‘fox’’, 220, backward], ...]
  2. 2.
    Words: Total Fixation Duration + Incoming/Outgoing Backward/Forward Saccades
    • Format: a list of word-level features: [word index, word, total fixation duration in ms, incoming forward saccades (from earlier word), incoming backward saccades (from later word), outgoing forward saccades (to later word), outgoing backward saccades (to earlier word)]
    • Example: [[4, ‘‘fox’’, 320, 2, 1, 3, 0], ...]
  3. 3.
    Words+Fixations
    • Format: two lists of word-level and fixation-level features: [fixated word index, fixated word, fixation duration in ms, direction of next saccade (backward / within / forward)] [word index, word, total fixation duration in ms, incoming forward saccades, incoming backward saccades, outgoing forward saccades, outgoing backward saccades]
    • Example: [[4, ‘‘fox’’, 220, backward], ...] [[4, ‘‘fox’’, 320, 2, 1, 3, 0], ...]
We further experiment with an alternative prompt that de-emphasizes the experimental setup and places more focus on the model’s generation task:
Task Description:
Your task is to generate the exact original question a reader had in mind before reading a given paragraph. The input data is (a/two) time series composed of (word-level/fixation-level/word-level and fixation-level) features.
Input Format:
You will receive the paragraph and eye movement data formatted as <FORMAT>
Expected Output:
Generate the exact original question provided to the reader, accurately inferred from the fixation patterns.
Instructions:
Produce only the exact original question as your output.
<PARAGRAPH>
<SCANPATH>
Appendix C Model Training and Hyperparameters
@@ C.1 Discriminative Models
For all neural models, we train with learning rates of $\{0.00001,0.00003,0.0001,0.0002\}$, following Shubi et al. (2024). Additionally, for all models that make use of word embeddings, we include both frozen and unfrozen language model variants in the search space.
  • For Haller RNN, we search over LSTM hidden sizes of $\{10,40,70,140\}$, using one layer and a dropout rate of $0.1$. The model is trained with and without freezing language model parameters.
@@
We train the deep-learning-based models for a maximum of 40 epochs, with early stopping after 8 epochs if no improvement in the validation error is observed. A single training epoch took roughly 5 minutes for RoBERTa-F, 10 minutes for Haller RNN, and 30 minutes for Llava and Llama. Each individual run was capped at 24 hours. Following Liu et al. (2019); Mosbach et al. (2021); Shubi et al. (2024), we use the AdamW optimizer (Loshchilov & Hutter, 2018) with a batch size of $16$. RoBERTa-F uses a linear warm-up ratio of $0.06$ and a weight decay of $0.1$. We standardize each eye movement feature using statistics computed on the training set, to zero mean unit variance.
Both reading-time informed embedding similarity baseline models use word embeddings from the RoBERTa-Large (Liu et al., 2019) language model.
C.2 Generative Models
@@
This model employs a teacher-forcing strategy. Specifically, loss computation is restricted exclusively to tokens corresponding to possible questions, contrasting the likelihood of generating the correct question against two distractors. The computed losses for each question candidate are inverted, serving as logits for softmax normalization and subsequently optimized via cross-entropy against the true label. The LLaVA backbone remains frozen, and training employs LoRA with dropout rate of 0.1, RS-LoRA, and rank $r=8$. The hyperparameter search includes learning rates of $1\times 10^{-5},3\times 10^{-5},1\times 10^{-4}$}, fixation embedding hidden sizes of 512,1024{512,1024}, and early stopping after 8 epochs without validation improvement.
All neural network-based models were trained using the PyTorch Lightning (Falcon & team, 2024) library on NVIDIA A100-40GB and L40S-48GB GPUs.
We utilize the Hugging Face implementation of LLaVA-OneVision, specifically LLaVA-hf/LLaVA-onevision-qwen2-0.5b-si-hf. Additionally, we employ the gpt-4o-2024-08-06 version of GPT-4o. We use Unsloth (version 2025.9.7) (Daniel Han & team, 2023) to fine-tune the Llama model. For the QA model we use LIAMF-USP/roberta-large-finetuned-race.